OpenStudy (ledah):

Multiplying radicals \[(4\sqrt{3})(-5\sqrt{21})\]

OpenStudy (anonymous):

\[\huge \color{blue}{(a \sqrt{b})(-c \sqrt{d}) = -ac \sqrt{b \times d}}\]

OpenStudy (ledah):

What is that exactly?

OpenStudy (anonymous):

put the values and solve..

OpenStudy (anonymous):

a = 4, b = 3. c = 5 and d= 21 Plug in the values..

OpenStudy (ledah):

Oh.

OpenStudy (ledah):

I don't know how to plug this into the formula i guess if that's what your going to call it

OpenStudy (anonymous):

You know the values of a, b, c and d??

OpenStudy (ledah):

yes.

OpenStudy (anonymous):

and you are getting that formula which I have written above??

OpenStudy (anonymous):

Substitute the value of a, b, c and d into that formula..

OpenStudy (anonymous):

\[\huge \color{blue}{(4 \sqrt{3})(-5 \sqrt{21}) = -4 \times 5 \sqrt{3 \times 21}}\]

OpenStudy (anonymous):

What is the problem you are facing in doing this much..??

OpenStudy (ledah):

I was never taught this in regular school. so I have to redo it over this summer in virtual school and my virtual school teacher is out for vacation until the 8th and thus begins my problem. I just don't get how to solve for this. It doesn't come easy to me seeing as I mix up numbers all the time

OpenStudy (anonymous):

What you get if you multiply -4 with 5??

OpenStudy (ledah):

-20